Hi. Thanks in advance. I can't get the solution to this. I have a work/rate word problem. Bill can do a job 20 minutes faster than Joe. If it takes both of them to do the whole job in 60 minutes, how long does it take Joe alone?
First, Joe's rate is 1/x and Bill's rate is 1/x-20.
What I came up with was this: 1 job = 1/x(60) + 1/x-20(60), which is really 1 job = 60/x + 60/x-20.
When I multiply both sides by the LCD, I get x(x-20) = 60x -120 + 60x.
Then I get x^2 + 20x = 120x - 120. After that, I know I goofed, cause I get x^2 = 140x -120, and the only way to solve that is to take the square root. I'm pretty sure the answer doesn't involve a square root.
Where do I go wrong?

Answer by stanbon(75887)

When I multiply both sides by the LCD, I get x(x-20) = 60x -120 + 60x.
Then I get x^2 + 20x = 120x - 120.
--------------
When you multiply both sides by the lcd you should get:
x^2-20x=....... NOT x^2+20x=.....
Cheers,
Stan H.